Verify that the given family of functions solves the differential equation.
(i) dy⁄dx= (1-2t)y², y=⅟(C – t + t²)
(ii) dy⁄dx= y² sin t, y=⅟(C + cos t)
1
Expert's answer
2020-06-29T18:45:00-0400

1) Given "y=\frac{1}{C – t + t²}" ________(1)

"\implies \frac{dy}{dt} = (-1)(C-t+t^2)^{-2} (-1+2t) = \frac{1-2t}{(C-t+t^2)^2} = (1-2t) y^2"

Hence, "y" given by (1) is solution of "\frac{dy}{dx} = (1-2t)y^2" .


2) Given "y=\frac{1}{C + cos t}" __________(2)

"\implies \frac{dy}{dt} = (-1)(C+cost)^{-2} (-sint) = y^2 sin(t)"

Hence, y given by (2) is solution of "\frac{dy}{dt} = y^2sin(t)"